OpenStudy (anonymous):

Simplify: 1 - (cos^2t/(1+sint))

OpenStudy (anonymous):

\[1-\frac{cos^2(t)}{1+sin(t)} = \frac{1+sin(t)}{1+sin(t)}-\frac{cos^2(t)}{1+sin(t)} = \frac{1+sin(t)+cos^2(t)}{1+sin(t)} \]

OpenStudy (anonymous):

And \[cos^2(t)=1-sin^2(t)\] So: \[\frac{1+sin(t)+cos^2(t)}{1+sin(t)} = \frac{1+sin(t)+1-sin^2(t)}{1+sin(t)} = \frac{-sin^2(t)+sin(t)+2}{1+sin(t)} = -sin(t)+2 \]